# references.bib

### v1

@@ 5 Ren–Xu counterexamples
In this way, we have $R_{m+1}=P_{k}^{3^{a}}.$ Moreover $Q_{k}(3,\mathbf{F}_{3})$ must be a non-counterexample by Lemma 5.4, so by our assumption $P_{k}$ is a generator. By Lemma 5.9, $P_{k}$ is not in $Q_{k+1}(3,\mathbf{F}_{3}),$ so the largest power of $(x_{1}-x_{2})$ dividing into $R_{m+1}$ must be $(x_{1}-x_{2})^{3^{a}(2k+1)}$ and $m+d=\frac{3^{a}(2k+1)-1}{2}$ by Lemma 4.1. Then for all $1\leq i\leq d,$
$\frac{2(m+i)+1-3^{a}(2k+1)}{2}$ $\leq\frac{2(m+d)+1-3^{a}(2k+1)}{2}$
$=0.$
Thus $R_{m+i}=P_{k}^{3^{a}}=R_{m+1}$ which is indeed in degree $33m+$ 3 by Lemma 5.8.
@@
We claim that for $d<i<2d$, $m+i\in X$. Let $I$ be the set of integers $h$ such that $P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ is a Ren–Xu counterexample for $Q_{h}(3,\mathbf{F}_{3})$. Notice that if $o\in I,$ then the set $\{s,s+1,s+2,\ldots,s+3^{a-1}-1\}$ containing $o$ and for which $s\equiv 3^{a-1}\pmod{3^{a}}$ gives exactly $I.$ We then note that $m+1\in I,$ yet $m\not\in I$ since $m\not\in X$. Thus $m\equiv 3^{a-1}-1\pmod{3^{a}}.$ Since $m+d=\frac{3^{a}(2k+1)-1}{2}\in I$ as well, we have $s=3^{a}k+3^{a-1},$ $m=3^{a}k+3^{a-1}-1,$ and $d=\frac{3^{a-1}+1}{2}.$ Then
$\frac{3^{a}(2k+1)-1}{2}\leq m+i<\frac{3^{a-1}+1}{2}+\frac{3^{a}(2k+1)-1}{2}=3^{a}k+2\cdot 3^{a-1},$
so $m+i\pmod{3^{a}}$ is in $\{3^{a-1},3^{a-1}+1,\ldots,2\cdot 3^{a-1}-1\}$ and thus in $X$.
@@
If $R_{m+i}=P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2b}$ where $b=\frac{2(m+i)+1-3^{a}(2k+1)}{2}$ for $d<i<2d,$ then $m+d=\frac{3^{a}(2k+1)-1}{2}$ implies $b=i-d.$ Thus $R_{m+i}=P_{k}^{3^{a}}\prod(x_{i}-x_{j})^{2(i-d)}$ has degree $3m+3+6(i-d)$ as desired.
@@
In [wang2022explicit], the first author proved that generators $Q_{m}(3,\mathbf{F}_{p})_{\mathrm{std}}$ for $p>3$ lie in $\mathbf{F}_{p}[x_{1}-x_{3},x_{2}-x_{3}]$ using that $\mathbf{F}_{p}[x_{1}-x_{3},x_{2}-x_{3},x_{1}+x_{2}+x_{3}]=\mathbf{F}_{p}[x_{1},x_{2},x_{3}].$ However this is not true for $p=3$ since $x_{1}-x_{3}+x_{2}-x_{3}=x_{1}+x_{2}+x_{3}$ in characteristic $3,$ so we instead consider the space $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}].$ From now on we say a polynomial’s degree in $x_{3}$ is with respect to the basis $\{x_{1}-x_{3},x_{2}-x_{3},x_{3}\}.$ Moreover, in [wang2022explicit] the first author defined the polynomial $M_{d}=(x_{1}+x_{2}-2x_{3})^{2\left\{\frac{d}{2}\right\}}(x_{1}-x_{3})^{\bigl{\lfloor}\frac{d}{2}\bigr{\rfloor}}(x_{2}-x_{3})^{\bigl{\lfloor}\frac{d}{2}\bigr{\rfloor}}$ for natural numbers $d$ and proved that homogeneous $s_{12}$-invariant elements of $\mathbf{F}_{p}[x_{1}-x_{3},x_{2}-x_{3}]/(x_{1}-x_{2})^{2}$ are equal to constant multiples of $M_{d}.$ Extending this gives that elements of $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}$ are polynomials in $x_{3}$ with coefficients that are constant multiples of $M_{d}.$ This gives us intuition for following lemmas.
@@
We prove the lemma by induction on $j.$
The base case for $n$ where $3\nmid n$ is $j=n-1$. If $j=n-1$ and $n\equiv 1\pmod{3},$ we can let $P=e_{3}^{\frac{n-1}{3}}e_{1}.$ If $n\equiv 2\pmod{3},$ we let $P=e_{3}^{\frac{n-2}{3}}e_{2}.$ The base case when $3|n$ is $j=n-2,$ so we can let $P=e_{1}e_{2}e_{3}^{\frac{n}{3}-1}$
Suppose that, when $3\nmid n,$ for all $i$ such that $n>i>j$ where $j\in\mathbf{N}$ and $10\leq j<n-$ 1 there exists a monomial in $e_{1},e_{2},e_{3}$ with degree $n$ and degree $i$ in $x_{3}.$ Suppose the same for when $3|n$ but with $n-1>i>j$ and $j<n-2$. Then there exists a monomial $m=e_{1}^{a}e_{2}^{b}e_{3}^{c}$ with degree $j+1$ in $x_{3}$ in $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}.$ If $b\neq 0$ we can take the monomial $e_{1}^{a+2}e_{2}^{b-1}e_{3}^{c}$ to be $P$ since it has degree $n$ and degree $j$ in $x_{3}$. If $b=0$ and $a,c>0,$ then we take $P=e_{1}^{a-1}e_{2}^{b+2}e_{3}^{c-1}.$ Finally we are left with the cases $a,b=0$ or $b,c=0.$ The former would imply $m=e_{3}^{\frac{n}{3}}$ is our monomial, but $3\nmid n,$ would imply $m$ is not a polynomial and $3|n$ implies $m$ has degree $j+1=n$ in $x_{3}$ and $j=n-1\not<n-2.$ For the latter case, we have that $a=n,$ so $m=e_{1}^{n}$ implies that $j+1=0$ which is below our range for $j$.
@@
For all $f_{i}\in\mathbf{F}_{3}$ and $n\not\equiv 0\pmod{3}$ there exists a $P\in\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$ such that
$P=f_{0}M_{n}x_{3}^{0}+f_{1}M_{n-1}x_{3}^{1}+\dotsm+f_{n-2}M_{2}x_{3}^{n-2}+f_{n-1}M_{1}x_{3}^{n-1}$
in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}.$ If $n\equiv 0\pmod{3},$ for all $f_{i}\in\mathbf{F}_{3}$ there exists a $P\in\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$ such that
$P=f_{0}M_{n}x_{3}^{0}+f_{1}M_{n-1}x_{3}^{1}+\dotsm+f_{n-2}M_{2}x_{3}^{n-2}$
@@
We prove this by induction.
For the base case when $n\not\equiv 0\pmod{3}$, we claim there exists coefficients $c_{i}\in\mathbf{F}_{3}$ such that the polynomial $c_{0}M_{n}x_{3}^{0}+c_{1}M_{n-1}x_{3}^{1}+\dotsm+c_{n-2}M_{2}x_{3}^{n-2}+c_{n-1}M_{1}x_{3}^{n-1}$ is in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}.$ The symmetric polynomial 00 satisfies these conditions and has degree 00 in $x_{3}.$ For the base case when $n\equiv 0\pmod{3},$ we claim there exists coefficients $c_{0},...,c_{n-2}$ such that the polynomial $c_{0}M_{n}x_{3}^{0}+c_{1}M_{n-1}x_{3}^{1}+\dotsm+c_{n-2}M_{2}x_{3}^{n-2}$ is in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}.$ The symmetric polynomial 00 satisfies this.
We consider the case where $n\not\equiv 0\pmod{3}.$ Suppose that for all $n\geq i>j$ there exists coefficients $c_{0},...,c_{i-1}$ such that for all $f_{i},f_{i+1},..,f_{n-1}$ there exists a symmetric polynomial $P$ such that
$P=c_{0}M_{n}x_{3}^{0}+c_{1}M_{n-1}x_{3}^{1}+\dotsm+c_{i-1}M_{n-i+1}x_{3}^{i-1}+f_{i}M_{n-i}x_{3}^{i}+\dotsm+f_{n-1}M_{1}x_{3}^{n-1}$
lies in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}$ where $j\in\mathbf{N}$ and $10\leq j\leq n-$ 1 . Moreover, suppose the polynomial $P$ exists such that it has degree in $x_{3}$ equal to the degree in $x_{3}$ in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}.$
Consider arbitrary coefficients $f_{j},f_{j+1},\ldots,f_{n-1}.$ If they are each $0,$ then we can take 00 to be our polynomial just like our base case. Otherwise, let $l$ be the greatest natural number $l\geq j$ such that $f_{l}\neq 0.$ If $l=j,$ by Lemma 5.13 there exists a monomial $m$ in $e_{1},e_{2},e_{3}$ with degree $j$ in $x_{3}$ and we may take $f_{j}m$ to be our symmetric polynomial.
@@
$S+(f_{j}-c_{j})m=c_{0}^{\prime}M_{n}x_{3}^{0}+c_{1}^{\prime}M_{n-1}x_{3}^{1}+\dotsm+c_{j-1}^{\prime}M_{n-j+1}x_{3}^{j-1}+f_{j}M_{n-j}x_{3}^{j}+\dotsm+f_{n-1}M_{1}x_{3}^{n-1}$
for some coefficients $c_{0}^{\prime},c_{1}^{\prime},\ldots,c_{j-1}^{\prime}.$ Moreover $S+(f_{j}-c_{j})m$ is still a symmetric polynomial and $m$ has degree $j$ in $x_{3}$ while $S$ has degree $l,$ so $S+(f_{j}-c_{j})m$ has degree $l$ as desired.
An identical argument holds for $n\equiv 0\pmod{3}.$
@@
Suppose that for all $i\leq m,$ if $i\not\in X$ then $Q_{i}(3,\mathbf{F}_{3})$ has a $13i+$ 1 degree generator where $m$ is a natural number. Then if $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $13m+$ 1, $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $63m+$ 6 .
@@
$\left((x_{1}+x_{2}+x_{3})\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)-x_{3}B\right)(x_{1}-x_{2})^{2m+1}$
@@
in degree $13m+$ 1 where $(x_{1}-x_{2})^{2m+1}(x_{1}+x_{2}-2x_{3})A^{\prime}$ and $(x_{1}-x_{2})^{2m+1}B^{\prime}$ are the generators of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}},$ $B$ is an $s_{12}$-invariant polynomial, and $\pi(A^{\prime})=\pi(B^{\prime})=B$.
For the greater degree generator, let $C=\left((x_{1}+x_{2}+x_{3})\pi\left(\frac{A^{\prime}-B^{\prime}}{3}\right)-x_{3}B\right)$. We would like to show there exists symmetric polynomials $P$ and $Q$ in degree $4$ and $5$ respectively such that
@@
This would then imply $(PC+QB)(x_{1}-x_{2})^{2m+1}\in Q_{m+1}(3,\mathbf{F}_{3})$ by Lemma 4.1. Consider writing

### v2

@@ 5 Ren–Xu counterexamples
In this way, we have $R_{m+1}=P_{k}^{3^{a}}$. Moreover, $Q_{k}(3,\mathbf{F}_{3})$ must be a non-counterexample by Lemma 5.4, so by our assumption $P_{k}$ is a generator. By Lemma 5.9, $P_{k}$ is not in $Q_{k+1}(3,\mathbf{F}_{3})$, so the largest power of $(x_{1}-x_{2})$ dividing into $R_{m+1}$ must be $(x_{1}-x_{2})^{3^{a}(2k+1)}$ and $m+d=\frac{3^{a}(2k+1)-1}{2}$ by Lemma 4.1. Then for all $1\leq j\leq d$,
$\frac{2(m+j)+1-3^{a}(2k+1)}{2}\leq\frac{2(m+d)+1-3^{a}(2k+1)}{2}=0.$
Thus $R_{m+j}=P_{k}^{3^{a}}=R_{m+1}$ which is indeed in degree $33m+$ 3 by Lemma 5.8.
We claim that for $d<j<2d$, $m+j\in X$. Let $I$ be the set of integers $h$ such that a Ren–Xu counterexample for $Q_{h}(3,\mathbf{F}_{3})$ is $P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2b}$ for some $b\in\mathbf{Z}_{\geq 0}$. By [11], $m\in I$ if and only if
$k+\frac{1}{3}\leq\frac{m}{3^{a}}\leq k+\frac{2}{3}-\frac{1}{3^{a}},$
@@
which implies $I$ is $\bigl{\{}s,s+1,s+2,\ldots,s+3^{a-1}-1\bigr{\}}$ for some $s\equiv 3^{a-1}\pmod{3^{a}}$. Then note that $m+1\in I$, yet $m\not\in I$ since $m\not\in X$. Thus $m\equiv 3^{a-1}-1\pmod{3^{a}}$. Since $m+d=\frac{3^{a}(2k+1)-1}{2}\in I$ as well, we have $s=3^{a}k+3^{a-1}$, $m=3^{a}k+3^{a-1}-1$, and $d=\frac{3^{a-1}+1}{2}$. Then
$\frac{3^{a}(2k+1)-1}{2}<m+j<\frac{3^{a-1}+1}{2}+\frac{3^{a}(2k+1)-1}{2}=3^{a}k+2\cdot 3^{a-1},$
so $m+j$ is in $I$ and thus in $X$.
@@
If $R_{m+j}=P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2b}$, where $b=\frac{2(m+j)+1-3^{a}(2k+1)}{2}$ for $d<j<2d$, then $m+d=\smash{\frac{3^{a}(2k+1)-1}{2}}$ implies $b=j-d$. Thus $R_{m+j}=P_{k}^{3^{a}}\prod(x_{i_{1}}-x_{i_{2}})^{2(j-d)}$ has degree $3m+3+6(j-d)$ as desired. ∎
@@
In [13], the first author proved that generators of $Q_{m}(3,\mathbf{F}_{p})_{\mathrm{std}}$ for $p>3$ lie in $\mathbf{F}_{p}[x_{1}-x_{3},\allowbreak{x_{2}-x_{3}}]$ using that $\mathbf{F}_{p}[x_{1}-x_{3},x_{2}-x_{3},x_{1}+x_{2}+x_{3}]=\mathbf{F}_{p}[x_{1},x_{2},x_{3}]$. However, this is not true for $p=3$ since $x_{1}-x_{3}+x_{2}-x_{3}=x_{1}+x_{2}+x_{3}$ in characteristic $3$, so we instead consider the space $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]$. From now on, we say a polynomial’s degree in $x_{3}$ is with respect to the basis $\{x_{1}-x_{3},x_{2}-x_{3},x_{3}\}$. Moreover, in [13] the first author defined the polynomial
$M_{d}=(x_{1}+x_{2}-2x_{3})^{2\{\frac{d}{2}\}}(x_{1}-x_{3})^{\lfloor\frac{d}{2}\rfloor}(x_{2}-x_{3})^{\lfloor\frac{d}{2}\rfloor}$
for natural numbers $d$ and proved that homogeneous $s_{12}$-invariant elements of $\mathbf{F}_{p}[x_{1}-x_{3},x_{2}-x_{3}]/(x_{1}-x_{2})^{2}$ are equal to constant multiples of $M_{d}$. Extending this gives that elements of $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}$ are polynomials in $x_{3}$ with coefficients that are constant multiples of $M_{d}$. Some further nice properties of $M_{d}$ are the following.
Lemma 5.13.
For any $j,j^{\prime}\in\mathbf{Z}_{\geq 0}$,
  1. $1$
    $(x_{1}+x_{2}+x_{3})M_{j}=M_{j+1}$ in $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}$.
  2. $2$
    $M_{j}M_{j^{\prime}}=M_{j+j^{\prime}}$ in $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}$.
Proof.
1. In $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}$, for $j\in Z_{\geq 0}$,
$(x_{1}+x_{2}+x_{3})M_{2j}$ $=(x_{1}+x_{2}+x_{3})(x_{1}-x_{3})^{j}(x_{2}-x_{3})^{j}=M_{2j+1}$
and
$(x_{1}+x_{2}+x_{3})M_{2j+1}$ $=(x_{1}+x_{2}+x_{3})^{2}(x_{1}-x_{3})^{j}(x_{2}-x_{3})^{j}$
$=(x_{1}-x_{3})(x_{2}-x_{3})M_{2j}=M_{2j+2}.$
2. From (1), we have $M_{j}=(x_{1}+x_{2}+x_{3})^{j}$ and $M_{j^{\prime}}=(x_{1}+x_{2}+x_{3})^{j^{\prime}}$ in $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}$, and our equality follows. ∎
This gives us intuition for the following lemmas.
@@ 1. In $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}$, for $j\in Z_{\geq 0}$,
We prove the lemma by decreasing induction on $j$.
The base case for $n$ where $3\nmid n$ is $j=n-1$. If $j=n-1$ and $n\equiv 1\pmod{3}$, we can let $P=e_{3}^{(n-1)/{3}}e_{1}$. If $n\equiv 2\pmod{3}$, we let $P=e_{3}^{(n-2)/{3}}e_{2}$. The base case when $3|n$ is $j=n-2$, so we can let $P=e_{1}e_{2}e_{3}^{(n/3)-1}$.
Suppose that, when $3\nmid n$, for all $j^{\prime}$ such that $n>j^{\prime}>j$ where $j\in\mathbf{N}$ and $10\leq j<n-$ 1 there exists a monomial in $e_{1}$, $e_{2}$, $e_{3}$ with degree $n$ and degree $j^{\prime}$ in $x_{3}$. Suppose the same for when $3|n$ but with $n-1>j^{\prime}>j$ and $j<n-2$. Then there exists a monomial $m=e_{1}^{a}e_{2}^{b}e_{3}^{c}$ with degree $j+1$ in $x_{3}$ in $\mathbf{F}_{3}[x_{1}-x_{3},x_{2}-x_{3},x_{3}]/(x_{1}-x_{2})^{2}$. If $b\neq 0$ we can take the monomial $e_{1}^{a+2}e_{2}^{b-1}e_{3}^{c}$ to be $P$ since it has degree $n$ and degree $j$ in $x_{3}$. If $b=0$ and $a,c>0$, then we take $P=e_{1}^{a-1}e_{2}^{b+2}e_{3}^{c-1}$. Finally, we are left with the cases $a,b=0$ or $b,c=0$. The former would imply $m=e_{3}^{\frac{n}{3}}$ is our monomial, but $3\nmid n$ would imply $m$ is not a polynomial and $3|n$ implies $m$ has degree $j+1=n$ in $x_{3}$ and $j=n-1\not<n-2$. For the latter case, we have that $a=n$, so ${m=e_{1}^{n}}$ implies that $j+1=0$ which is below our range for $j$. ∎
@@
For all $f_{j}\in\mathbf{F}_{3}$ and $n\not\equiv 0\pmod{3}$, there exists a $P\in\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$ such that
$P=f_{0}M_{n}x_{3}^{0}+f_{1}M_{n-1}x_{3}^{1}+\dots+f_{n-2}M_{2}x_{3}^{n-2}+f_{n-1}M_{1}x_{3}^{n-1}$
in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}$. If $n\equiv 0\pmod{3}$, for all $f_{j}\in\mathbf{F}_{3}$ there exists a $P\in\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}$ such that
$P=f_{0}M_{n}x_{3}^{0}+f_{1}M_{n-1}x_{3}^{1}+\dots+f_{n-2}M_{2}x_{3}^{n-2}$
@@
A weaker statement is that there exists some fixed $c_{0},c_{1},\dots,c_{j}\in\mathbf{F}_{3}$ such that for all $f_{j+1},f_{j+2},\dots,f_{n-1}\in\mathbf{F}_{3}$, there exists a symmetric polynomial
$P$ $\equiv c_{0}M_{n}x_{3}^{0}+c_{1}M_{n-1}x_{3}^{1}+\dots+c_{j}M_{n-j}x_{3}^{j}$
$\quad+f_{j+1}M_{2}x_{3}^{n-j-1}+f_{j+2}M_{1}x_{3}^{n-j-2}+\dots+f_{n-1}M_{1}x_{3}^{n-1}\pmod{(x_{1}-x_{2})^{2}},$
when $n\not\equiv 3\pmod{3}$ and $j\in\mathbf{Z}_{\geq 0}$. A similar weaker statement can be made for the $n\equiv 0\pmod{3}$ case. We prove the statement in the lemma by induction on this $j$.
For the base case when $n\not\equiv 0\pmod{3}$, we claim there exists coefficients $c_{j}\in\mathbf{F}_{3}$ such that the polynomial $c_{0}M_{n}x_{3}^{0}+c_{1}M_{n-1}x_{3}^{1}+\dots+c_{n-2}M_{2}x_{3}^{n-2}+c_{n-1}M_{1}x_{3}^{n-1}$ is in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}$. The symmetric polynomial 00 satisfies these conditions and has degree 00 in $x_{3}$. For the base case when $n\equiv 0\pmod{3}$, we claim there exists coefficients $c_{0},\dots,c_{n-2}$ such that the polynomial $c_{0}M_{n}x_{3}^{0}+c_{1}M_{n-1}x_{3}^{1}+\dots+c_{n-2}M_{2}x_{3}^{n-2}$ is in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}$. The symmetric polynomial 00 satisfies this.
We consider the case where $n\not\equiv 0\pmod{3}$. Suppose that for all $n\geq j^{\prime}>j$ there exists coefficients $c_{0},\dots,c_{j^{\prime}-1}$ such that for all $f_{j^{\prime}},f_{j^{\prime}+1},\dots,f_{n-1}$ there exists a symmetric polynomial $P$ such that
$P=c_{0}M_{n}x_{3}^{0}+c_{1}M_{n-1}x_{3}^{1}+\dots+c_{j^{\prime}-1}M_{n-j^{\prime}+1}x_{3}^{j^{\prime}-1}+f_{j^{\prime}}M_{n-j^{\prime}}x_{3}^{j^{\prime}}+\dots+f_{n-1}M_{1}x_{3}^{n-1}$
lies in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}$, where $j\in\mathbf{N}$, $10\leq j\leq n-$ 1. Moreover, suppose the polynomial $P$ exists such that it has degree in $x_{3}$ equal to the degree in $x_{3}$ in $\mathbf{F}_{3}[x_{1},x_{2},x_{3}]^{S_{3}}/(x_{1}-x_{2})^{2}$.
Consider arbitrary coefficients $f_{j},f_{j+1},\ldots,f_{n-1}$. If they are each 00, then we can take 00 to be our polynomial just like our base case. Otherwise, let $l$ be the greatest natural number $l\geq j$ such that $f_{l}\neq 0$. If $l=j$, by Lemma 5.14 there exists a monomial $m$ in $e_{1}$, $e_{2}$, $e_{3}$ with degree $j$ in $x_{3}$ and we may take $f_{j}m$ to be our symmetric polynomial.
@@
$S+(f_{j}-c_{j})m=c_{0}^{\prime}M_{n}x_{3}^{0}+c_{1}^{\prime}M_{n-1}x_{3}^{1}+\dots+c_{j-1}^{\prime}M_{n-j+1}x_{3}^{j-1}$
${}+f_{j}M_{n-j}x_{3}^{j}+\dots+f_{n-1}M_{1}x_{3}^{n-1}$
for some coefficients $c_{0}^{\prime},c_{1}^{\prime},\ldots,c_{j-1}^{\prime}$. Moreover, $S+(f_{j}-c_{j})m$ is still a symmetric polynomial and $m$ has degree $j$ in $x_{3}$ while $S$ has degree $l$, so $S+(f_{j}-c_{j})m$ has degree $l$ as desired.
An identical argument holds for $n\equiv 0\pmod{3}$. ∎
@@
Suppose that for all $k\leq m$, if $k\not\in X$ then $Q_{k}(3,\mathbf{F}_{3})$ has a $13k+$ 1 degree generator, where $m$ is a natural number. Then if $Q_{m}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree ${3m+1}$, $Q_{m+1}(3,\mathbf{F}_{3})_{\mathrm{sign-triv}}$ has a generator in degree $63m+$ 6.
@@
$\biggl{(}(x_{1}+x_{2}+x_{3})\pi\biggl{(}\frac{A^{\prime}-B^{\prime}}{3}\biggr{)}-x_{3}B\biggr{)}(x_{1}-x_{2})^{2m+1}$
@@
in degree $13m+$ 1, where $(x_{1}-x_{2})^{2m+1}(x_{1}+x_{2}-2x_{3})A^{\prime}$ and $(x_{1}-x_{2})^{2m+1}B^{\prime}$ are the generators of $Q_{m}(3,\mathbf{Q})_{\mathrm{std}}$, $B$ is an $s_{12}$-invariant polynomial, and $\pi(A^{\prime})=\pi(B^{\prime})=B$.
For the greater degree generator, let $C=\big{(}(x_{1}+x_{2}+x_{3})\pi\big{(}\frac{A^{\prime}-B^{\prime}}{3}\big{)}-x_{3}B\big{)}$. We would like to show there exists symmetric polynomials $P$ and $Q$ in degree $4$ and $5$ respectively such that
@@
Since $\frac{PC+BQ}{(x_{1}-x_{2})^{2}}$ is still $s_{12}$-invariant, this would then imply $(PC+QB)(x_{1}-x_{2})^{2m+1}\in Q_{m+1}(3,\mathbf{F}_{3})$ by Lemma 4.1. Consider writing